### v1

# CONDA 2024 Shared Task on Evidence of Data Contamination: First Compilation Report

Abstract
The 1st Workshop on Data Contamination (CONDA 2024) focuses on all relevant aspects of data contamination in natural language processing, where data contamination is understood as situations where evaluation data is included in pre-training corpora used to train large scale models, compromising evaluation results. The workshop fostered a shared task to collect evidence on data contamination in current available datasets and models. The goal of the shared task and associated database is to assist the community in understanding the extent of the problem and to assist researchers in avoiding reporting evaluation results on known contaminated resources. The shared task provides a structured, centralized public database for the collection of contamination evidence, open to contributions from the community via GitHub pool requests. This first compilation paper is based on 566 reported entries over 91 contaminated sources from a total of 23 contributors. The details of the individual contamination events are available in the platform.1 The platform continues to be online, open to contributions from the community.
Data Contamination Report from the 2024 CONDA Shared Task
Oscar Sainz1 Iker García-Ferrero1 Alon Jacovi2 Jon Ander Campos3Yanai Elazar4,5Eneko Agirre1Yoav Goldberg2,4 Wei-Lin Chen6,7Jenny Chim8Leshem Choshen9,10Luca D’Amico-Wong11 Melissa Dell11Run-Ze Fan12Shahriar Golchin13Yucheng Li14Pengfei Liu12 Bhavish Pahwa15Ameya Prabhu17Suryansh Sharma18Emily Silcock11 Kateryna Solonko15David Stap19Mihai Surdeanu20Yu-Min Tseng21 Vishaal Udandarao16,17,22Zengzhi Wang12Ruijie Xu12Jinglin Yang11 1HiTZ Center - Ixa, University of the Basque Country UPV/EHU  2Bar Ilan University  3Cohere 4Allen Institute for Artificial Intelligence  5University of Washington  6National Taiwan University 7University of Virginia  8Queen Mary University of London  9MIT-IBM Watson AI Lab  10MIT 11Harvard University  12Shanghai Jiao Tong University  13University of Arizona  14University of Surrey 15Microsoft Research  16Tübingen AI Center  17University of Tübingen 18Indian Institute of Technology Kharagpur  19University of Amsterdam  20University of Arizona 21Microsoft  22University of Cambridge Contact: conda-workshop@googlegroups.com
1 Introduction
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=3cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [Contamination detection methods, fill=gray!45, parent [Data-based, for tree= data_based [Proprietary data, data_based [GPT-3 Brown et al. (2020), style = data_based_work] [FLAN Wei et al. (2022), style = data_based_work] [GLaM Du et al. (2022a), style = data_based_work] [PaLM Chowdhery et al. (2022), style = data_based_work] [PaLM-2 Anil et al. (2023), style = data_based_work] [GPT-4 OpenAI et al. (2024), style = data_based_work] ] [Open data, data_based [ Dodge et al. (2021), style = data_based_work] [ Silcock et al. (2023), style = data_based_work] [ Muennighoff et al. (2023), style = data_based_work] [ Azerbayev et al. (2023), style = data_based_work] [ Elazar et al. (2024), style = data_based_work] [ Riddell et al. (2024), style = data_based_work] [ Li et al. (2024), style = data_based_work] ] ] [Model-based, for tree=model_based [Closed model, model_based [ Sainz et al. (2023b), style = model_based_work] [ Golchin and Surdeanu (2024b), style = model_based_work] [ Golchin and Surdeanu (2024a), style = model_based_work] [ Dong et al. (2024), style = model_based_work] [ Ranaldi et al. (2024), style = model_based_work] [ Enis and Hopkins (2024), style = model_based_work] ] [Open model, model_based [ Deng et al. (2024), style = model_based_work] [ Oren et al. (2024), style = model_based_work] [ Xu et al. (2024), style = model_based_work] ] ] ]
Figure 1: Taxonomy of papers that report contamination evidence. Including LLM’s papers and technical reports, papers about methods for detecting contamination, and papers about corpus analysis.
Data contamination, where evaluation data is inadvertently included in pre-training corpora of large-scale models, and language models (LMs) in particular, has become a concern in recent times Sainz et al. (2023a); Jacovi et al. (2023). The growing scale of both models and data, coupled with massive web crawling, has led to the inclusion of segments from evaluation benchmarks in the pre-training data of LMs Dodge et al. (2021); OpenAI et al. (2024); Anil et al. (2023); Elazar et al. (2024). The scale of internet data makes it difficult to prevent this contamination from happening, or even detect when it has happened Bommasani et al. (2022); Mitchell et al. (2023).
Crucially, when evaluation data becomes part of pre-training data, it introduces biases and can artificially inflate the performance of LMs on specific tasks or benchmarks (Magar and Schwartz, 2022; Magnusson et al., 2023; Merrill et al., 2024). This poses a challenge for fair and unbiased evaluation of models, as their performance may not accurately reflect their generalization capabilities (Hupkes et al., 2023). And similarly to pre-training contamination, the contamination can also occur during the fine-tuning stage even after a model has been deployed as an API Balloccu et al. (2024).
Although a growing number of papers and state-of-the-art models mention issues of data contamination Brown et al. (2020); Wei et al. (2022); Chowdhery et al. (2022); OpenAI et al. (2024); Anil et al. (2023); Touvron et al. (2023), there is little in the way of organized and compiled knowledge about real, documented cases of contamination in practice Sainz et al. (2023a). Addressing data contamination is a shared responsibility among researchers, developers, and the broader community.
This report compiles the evidence reported in the Data Contamination Database [^1] https://huggingface.co/spaces/CONDA-Workshop/Data-Contamination-Database as part of the Data Contamination Workshop. [^2] https://conda-workshop.github.io/ As the Shared Task of the workshop, researchers were invited to discover cases of contamination in available corpora and models, and submit evidence of their discovery. The submissions to the database were collected and compiled on June 23rd, 2024, to be included in this report, but the database continues to run and grow. Overall we collected 566 submissions from 23 contributors, where each submission included a detailed contamination report, indicating the estimated percentage of contaminated data. We continue to operate the database, and expect to update it with newer datasets and models as they come out, as well as new report about existing contaminated (or uncontaminated) evaluations.
This report first presents the methodology for collecting evidence, as well as existing papers that report data contamination (Section 2). We also report the evidence collected in the Data Contamination Database (Section 3), followed by an overview of the trends and statistics in the database, that inform a high-level perspective on the state of data contamination in NLP today (Section 4).
2 Methodology and Previous Work
Figure 2: Number of test sets reported for each corpus often used in pre-training.
Figure 3: Number of test sets reported for each pre-trained model.
Collecting all the contamination evidence —or lack of it— was done openly, through pull requests, and subject to discussions before the admission. Contributors were asked to fill in the information about several aspects, such as the contaminated resource (a training corpus or model), the evaluation dataset which was found in the contaminated source, a breakdown of the percentage of contamination found in each split of the dataset (train, development, and test), an optional reference to a paper that describes the methodology behind the submission, as well as whether the contamination detection method was data-based or model-based. The contributions provided the HuggingFace Hub id of models, corpus, and datasets when possible. In addition, contributors must provide the evidence or a reference to the scientific paper that reported the evidence originally. Figure 1 shows the taxonomy of the papers that reported contamination evidence in the shared task. [^3] Note that there are many other works on data contamination detection. In this report we focus on works that were used to detect contamination for this report. We leave a more detailed coverage survey for future work. We split the these methods into two: data-based and model-based approaches.
Data-based approaches.
are methods that inspect the pre-training corpora to find contamination evidence. Data-based approaches typically involve string or sub-string matching techniques such as 13-gram overlap Brown et al. (2020); Wei et al. (2022), 50-character overlap OpenAI et al. (2024) or even full-string overlap Elazar et al. (2024). In Figure 1 we differentiate between Proprietary and Open data. Papers that fall in the category of Proprietary data are usually LLMs technical reports that run post-hoc data contamination evaluations to identify and remove evaluation instances that appear in the pre-training corpora Brown et al. (2020); Wei et al. (2022); OpenAI et al. (2024). Papers that fall in the open data category usually involve corpus analysis tools Dodge et al. (2021); Elazar et al. (2024) or LLMs with publicly available pre-training data Azerbayev et al. (2023).
Model-based approaches.
are those methods that try to estimate the contamination of a model by prompting or analyzing the output, without accessing the pre-training data. These methods are formulated as Membership Inference Attacks (MIA) and range from asking LLMs to generate verbatim of the actual evaluation data Sainz et al. (2023b); Golchin and Surdeanu (2024b) to analyzing the actual output probabilities given by the model Oren et al. (2024). We differentiate between methods applicable to closed and open models. Methods applicable to closed models are usually applicable to open models, but not the other way around due to the limitations established by the API or interface providers.
The collected evidences come from different approaches and sources, making them hardly comparable. For transparency, we included in the database information about the source of the evidence and the link to the discussion. We encourage the users to assess how the evidence was collected for their datasets of interest.
3 Compilation of Evidence
The report includes 42 contaminated sources (training corpora or models), 91 datasets, and 566 contamination entries, including 432 contamination events (20 train-set, 95 dev-set, 317 test-set) and 144 non-contamination events, where a contamination event is taken as any report above 0% of contamination. The database contains, for each split (train, dev, and test) of each evaluation dataset, what percentage was found to be contaminated by a subset of the contamination sources (corpora or models). We analyze separately the contaminated corpora and models.
Figure 4: Percentage of contaminated report per task
Contaminated corpora.
Figure 3 shows the number of reported test sets for each corpus often used to pre-train language models. The reported corpora are mainly based on CommonCrawl snapshots, GitHub, or a mix of sources. For CommonCrawl-based corpora, there are 35 events reported for C4 Raffel et al. (2023), 32 for RedPajama v2 Computer (2023), 29 for OSCAR Jansen et al. (2022); Abadji et al. (2022, 2021); Kreutzer et al. (2022); Ortiz Su’arez et al. (2020); Ortiz Su’arez et al. (2019) and 6 for CommonCrawl Rana (2010) itself. Regarding the GitHub data, there are 2 events reported for the TheStack Kocetkov et al. (2022) project. The corpora with various sources, the Pile Gao et al. (2020) and ProofPile Azerbayev et al. (2023), have 30 and 2 reported contamination events respectively. There is also 1 report for xP3 Muennighoff et al. (2022), which is a collection of prompts for different NLP datasets. [^4] The report indicates the use of validation data from a specific dataset as training.
Table 1 shows for each corpus often used to pre-train language models, the contamination events involving development or test splits. Please refer to the online database for full details of each report.
Contaminated Source Evaluation Set
allenai/c4 Raffel et al. (2023), sem_eval_2014_task_1 Marelli et al. (2014), race, nyu-mll/glue Wang et al. (2019b), amazon_reviews_multi Keung et al. (2020), liar Wang (2017), reddit_tifu Kim et al. (2018), stsb_multi_mt May (2021), wiki_qa Yang et al. (2015), gigaword Graff et al. (2003), piqa Bisk et al. (2020), esnli Camburu et al. (2018), scitail Khot et al. (2018), snli Bowman et al. (2015), ibm/duorc Saha et al. (2018), math_qa Amini et al. (2019), swag Zellers et al. (2018), wiki_bio Lebret et al. (2016), xnli Conneau et al. (2018), allenai/scicite Cohan et al. (2019), aeslc Zhang and Tetreault (2019), billsum Kornilova and Eidelman (2019), AMR-to-Text, winograd_wsc Levesque et al. (2012), squadshifts Miller et al. (2020), head_qa Vilares and Gómez-Rodríguez (2019), xsum Narayan et al. (2018), health_fact Kotonya and Toni (2020), EdinburghNLP/xsum Narayan et al. (2018), UCLNLP/adversarial_qa Bartolo et al. (2020), paws Zhang et al. (2019), sick, super_glue Wang et al. (2019a), paws-x Yang et al. (2019), scan, lama Petroni et al. (2019, 2020)
CommonCrawl Rana (2010) allenai/ai2_arc Clark et al. (2018), tau/commonsense_qa Talmor et al. (2019), ceval/ceval-exam Huang et al. (2023), cais/mmlu Hendrycks et al. (2021a), Rowan/hellaswag Zellers et al. (2019), winogrande Levesque et al. (2012)
EleutherAI/pile Gao et al. (2020) sem_eval_2014_task_1 Marelli et al. (2014), nyu-mll/glue Wang et al. (2019b), amazon_reviews_multi Keung et al. (2020), mbpp, openai_humaneval Chen et al. (2021), liar Wang (2017), stsb_multi_mt May (2021), wiki_qa Yang et al. (2015), gigaword Graff et al. (2003), piqa Bisk et al. (2020), esnli Camburu et al. (2018), scitail Khot et al. (2018), snli Bowman et al. (2015), ibm/duorc Saha et al. (2018), swag Zellers et al. (2018), xnli Conneau et al. (2018), allenai/scicite Cohan et al. (2019), aeslc Zhang and Tetreault (2019), billsum Kornilova and Eidelman (2019), winograd_wsc Levesque et al. (2012), squadshifts Miller et al. (2020), head_qa Vilares and Gómez-Rodríguez (2019), xsum Narayan et al. (2018), health_fact Kotonya and Toni (2020), UCLNLP/adversarial_qa Bartolo et al. (2020), paws Zhang et al. (2019), sick, super_glue Wang et al. (2019a), paws-x Yang et al. (2019), scan
oscar-corpus/OSCAR-2301 Jansen et al. (2022); Abadji et al. (2022, 2021); Kreutzer et al. (2022); Ortiz Su’arez et al. (2020); Ortiz Su’arez et al. (2019) sem_eval_2014_task_1 Marelli et al. (2014), crows_pairs Nangia et al. (2020), nyu-mll/glue Wang et al. (2019b), race, amazon_reviews_multi Keung et al. (2020), openai_humaneval Chen et al. (2021), liar Wang (2017), stsb_multi_mt May (2021), wiki_qa Yang et al. (2015), gigaword Graff et al. (2003), piqa Bisk et al. (2020), esnli Camburu et al. (2018), scitail Khot et al. (2018), snli Bowman et al. (2015), math_qa Amini et al. (2019), swag Zellers et al. (2018), xnli Conneau et al. (2018), allenai/scicite Cohan et al. (2019), aeslc Zhang and Tetreault (2019), billsum Kornilova and Eidelman (2019), winograd_wsc Levesque et al. (2012), squadshifts Miller et al. (2020), head_qa Vilares and Gómez-Rodríguez (2019), xsum Narayan et al. (2018), health_fact Kotonya and Toni (2020), UCLNLP/adversarial_qa Bartolo et al. (2020), paws Zhang et al. (2019), sick, super_glue Wang et al. (2019a)
togethercomputer/RedPajama-Data-V2 Computer (2023) sem_eval_2014_task_1 Marelli et al. (2014), race, nyu-mll/glue Wang et al. (2019b), amazon_reviews_multi Keung et al. (2020), liar Wang (2017), stsb_multi_mt May (2021), wiki_qa Yang et al. (2015), gigaword Graff et al. (2003), piqa Bisk et al. (2020), esnli Camburu et al. (2018), scitail Khot et al. (2018), snli Bowman et al. (2015), ibm/duorc Saha et al. (2018), math_qa Amini et al. (2019), swag Zellers et al. (2018), xnli Conneau et al. (2018), allenai/scicite Cohan et al. (2019), aeslc Zhang and Tetreault (2019), billsum Kornilova and Eidelman (2019), winograd_wsc Levesque et al. (2012), squadshifts Miller et al. (2020), head_qa Vilares and Gómez-Rodríguez (2019), xsum Narayan et al. (2018), health_fact Kotonya and Toni (2020), UCLNLP/adversarial_qa Bartolo et al. (2020), mc_taco, paws Zhang et al. (2019), samsum Gliwa et al. (2019), sick, super_glue Wang et al. (2019a), paws-x Yang et al. (2019), scan
bigscience/xP3 Muennighoff et al. (2022) facebook/flores NLLB-Team et al. (2022)
EleutherAI/proof-pile-2 Azerbayev et al. (2023) gsm8k Cobbe et al. (2021), hendrycks/competition_math Hendrycks et al. (2021b)
bigcode/the-stack Kocetkov et al. (2022) openai_humaneval Chen et al. (2021), mbpp
Table 1: A summary of the dev or test sets found at above 0% contamination in each corpus often used to pre-train models.
Contaminated Source Evaluation Set
GPT-3 Brown et al. (2020) Reversed Words , race , quac Choi et al. (2018), Anagrams 1 , Cycled Letters , mandarjoshi/trivia_qa Joshi et al. (2017), ibragim-bad/arc_easy Clark et al. (2018), SAT Analogies, piqa Bisk et al. (2020), Rowan/hellaswag Zellers et al. (2019), wmt/wmt16 Bojar et al. (2016), stanfordnlp/coqa Reddy et al. (2019), cimec/lambada Paperno et al. (2016), natural_questions Kwiatkowski et al. (2019), winograd_wsc Levesque et al. (2012), ucinlp/drop Dua et al. (2019), rmanluo/RoG-webqsp, rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016), allenai/openbookqa Mihaylov et al. (2018), Symbol Insertion, Anagrams 2 , super_glue Wang et al. (2019a), ibragim-bad/arc_challenge Clark et al. (2018), facebook/anli Nie et al. (2020)
GPT-3.5 Brown et al. (2020) samsum Gliwa et al. (2019), yelp_review_full Zhang et al. (2015a), imdb Maas et al. (2011), ag_news Zhang et al. (2015b), nyu-mll/glue Wang et al. (2019b), conll2003 Tjong Kim Sang and De Meulder (2003), winogrande Levesque et al. (2012), rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016), cais/mmlu Hendrycks et al. (2021a), EdinburghNLP/xsum Narayan et al. (2018), allenai/openbookqa Mihaylov et al. (2018), xlangai/spider Yu et al. (2018), truthful_qa Lin et al. (2022)
GPT-4 OpenAI et al. (2024) samsum Gliwa et al. (2019), yelp_review_full Zhang et al. (2015a), gsm8k Cobbe et al. (2021), imdb Maas et al. (2011), ibragim-bad/arc_challenge Clark et al. (2018), nyu-mll/glue Wang et al. (2019b), ucinlp/drop Dua et al. (2019), winogrande Levesque et al. (2012), openai_humaneval Chen et al. (2021), ag_news Zhang et al. (2015b), EdinburghNLP/xsum Narayan et al. (2018), cais/mmlu Hendrycks et al. (2021a), Rowan/hellaswag Zellers et al. (2019), allenai/openbookqa Mihaylov et al. (2018), truthful_qa Lin et al. (2022), bigbench Srivastava et al. (2023)
PaLM 2 Anil et al. (2023) EdinburghNLP/xsum Narayan et al. (2018), csebuetnlp/xlsum Hasan et al. (2021), wiki_lingua Ladhak et al. (2020)
GPT-3.5-turbo Brown et al. (2020) openai_humaneval Chen et al. (2021), HumanEval_R Chen et al. (2021)
FLAN Wei et al. (2022) natural_questions Kwiatkowski et al. (2019), mandarjoshi/trivia_qa Joshi et al. (2017), story_cloze Sharma et al. (2018), piqa Bisk et al. (2020), super_glue Wang et al. (2019a), ibragim-bad/arc_challenge Clark et al. (2018), ucinlp/drop Dua et al. (2019), rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016), ibragim-bad/arc_easy Clark et al. (2018), Rowan/hellaswag Zellers et al. (2019), allenai/openbookqa Mihaylov et al. (2018), facebook/anli Nie et al. (2020), winogrande Levesque et al. (2012), wmt/wmt16 Bojar et al. (2016)
GLaM Du et al. (2022a) stanfordnlp/coqa Reddy et al. (2019), natural_questions Kwiatkowski et al. (2019), mandarjoshi/trivia_qa Joshi et al. (2017), story_cloze Sharma et al. (2018), cimec/lambada Paperno et al. (2016), piqa Bisk et al. (2020), super_glue Wang et al. (2019a), ibragim-bad/arc_challenge Clark et al. (2018), race , quac Choi et al. (2018), winograd_wsc Levesque et al. (2012), rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016), ibragim-bad/arc_easy Clark et al. (2018), Rowan/hellaswag Zellers et al. (2019), allenai/openbookqa Mihaylov et al. (2018), facebook/anli Nie et al. (2020), winogrande Levesque et al. (2012)
LLaMa 2-13B Touvron et al. (2023) allenai/openbookqa Mihaylov et al. (2018), winogrande Levesque et al. (2012), truthful_qa Lin et al. (2022)
Mistral-7B Jiang et al. (2023) allenai/openbookqa Mihaylov et al. (2018), winogrande Levesque et al. (2012), truthful_qa Lin et al. (2022), cais/mmlu Hendrycks et al. (2021a)
PaLM Chowdhery et al. (2022) cimec/lambada Paperno et al. (2016), super_glue Wang et al. (2019a), ibragim-bad/arc_challenge Clark et al. (2018), winograd_wsc Levesque et al. (2012), rmanluo/RoG-webqsp, rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016), mandarjoshi/trivia_qa Joshi et al. (2017), ibragim-bad/arc_easy Clark et al. (2018)
Claude 3 Opus facebook/flores NLLB-Team et al. (2022)
bigscience/bloomz Muennighoff et al. (2022) facebook/flores NLLB-Team et al. (2022)
bigscience/mt0-* Muennighoff et al. (2022) facebook/flores NLLB-Team et al. (2022)
BAAI/Aquila2-34B gsm8k Cobbe et al. (2021), hendrycks/competition_math Hendrycks et al. (2021b)
BAAI/AquilaChat2-34B gsm8k Cobbe et al. (2021)
EleutherAI/llemma_* Azerbayev et al. (2023) gsm8k Cobbe et al. (2021), hendrycks/competition_math Hendrycks et al. (2021b)
Qwen/Qwen-1_8B Bai et al. (2023) gsm8k Cobbe et al. (2021), hendrycks/competition_math Hendrycks et al. (2021b)
BAAI/Aquila2-7B hendrycks/competition_math Hendrycks et al. (2021b)
Qwen/Qwen-* Bai et al. (2023) hendrycks/competition_math Hendrycks et al. (2021b)
THUDM/chatglm3-6b Du et al. (2022b) hendrycks/competition_math Hendrycks et al. (2021b)
internlm/internlm2-* Cai et al. (2024) hendrycks/competition_math Hendrycks et al. (2021b)
mistralai/Mistral-7B-v0.1 Jiang et al. (2023) ibragim-bad/arc_easy Clark et al. (2018)
Table 2: A summary of the dev or test sets found at above 0% contamination in each reported model. The "*" is used to indicate the different versions or sizes of the models.
Contaminated models.
Figure 3 details the number of contamination events involving test sets that were reported, organised according to each pre-trained model. Most reported evidence is for closed models, for instance: 24 for GPT-3 Brown et al. (2020), 17 for GLaM Du et al. (2022a), 16 for GPT-4 OpenAI et al. (2024), 13 for GPT-3.5 Brown et al. (2020), 8 for PaLM Chowdhery et al. (2022), 3 for PaLM-2 Anil et al. (2023), 2 for GPT-3.5 Turbo Brown et al. (2020) and 1 for Calude 3 Opus. In the case of open models: there are 14 reported events for models fine-tuned with FLAN data Wei et al. (2022), 5 for Mistral Jiang et al. (2023), 3 for Llama 2 Touvron et al. (2023), 2 for Qwen Bai et al. (2023), Llema Azerbayev et al. (2023) and Aquila 2; and a single one for mT0 and Bloom-Z Muennighoff et al. (2022).
Table 2 shows for each pre-trained language model, the contamination events involving development or test splits. Please refer to the online database for full details of each report.
4 Analysis of the Reported Data
In this section, we analyze the reported entries to understand the report’s data better.
Figure 5: Number of downloads in the HuggingFace hub of the datasets in the report.
Reported tasks.
Figure 4 shows the percentage of data contamination per task. We use the task_id assigned to each dataset in the Hugging Face hub. Text-scoring, QA, and multiple-choice-qa are among the most contaminated task types. Figure  5 shows the number of downloads for every dataset in the report. We measure the total number of downloads from the Hugging Face hub. [^5] https://huggingface.co/docs/datasets Since one model may be reported as contaminated with a dataset while another model may not, we have entries of both being compromised and non-compromised for some datasets. Relating both tables, we can see that the tasks reported as the most contaminated include very popular datasets such as MMLU (multiple-choice-qa), GLUE (text-scoring), and ai2_arc (multiple-choice-qa), which are standard benchmarks for measuring the performance of LLMs. These benchmarks, as well as other very popular benchmarks reported in instances of data contamination, such as hellaswag or gsm8k are implemented in community leaderboards such as the Open LLM Leaderboard. [^6] https://hf.co/spaces/open-llm-leaderboard/
Figure 6: Year of publication of the contaminated test sets reported for each model.
Figure 7: Publication year of the test sets included in the data contamination report.
Year of publication of the reported data.
Figure 7 shows the percentage of total test sets included in contamination events per year. We present data for test sets in both contamination events (>0% contamination) and non-contamination events (0% contamination). Most of the reported datasets correspond to the 2018 to 2021 period.
We further explore the relationship between the year of publication of the datasets and instances of contamination by examining the reported data contamination for the three models with the most instances of data contamination: GPT-4, GPT-3, and GPT-3.5. As expected based on the models’ release dates, Figure 7 shows that more recently released models are contaminated with more recently released datasets. For instance, GPT-3, launched in 2020, is predominantly contaminated with datasets from 2016, while GPT-4, released in 2023, is mainly contaminated with datasets from 2018 to 2022.
5 Conclusions
Data contamination has become a significant concern in recent times. Consequently, a growing number of papers and state-of-the-art models mention issues of data contamination. In the CONDA 2024 Shared Task on Evidence of Data Contamination, we have collected and compiled a comprehensive database of available evidence on data contamination in currently available datasets and models. This report includes 566 contamination entries over 91 contaminated sources from a total of 23 contributors. With this shared task, we provide a structured, centralized platform for contamination evidence collection to help the community understand the extent of the problem and to assist researchers in avoiding reporting evaluation results on known contaminated resources. Given the large exploration space, this report does not cover all cases, but a small sample that were reported during our shared task period, in the midst of 2024. We welcome further submissions to the database, and plan to keep this database up-to-date as it provides a valuable source of information for the research community.
Acknowledgments
We are grateful to Hugging Face and Clémentine Fourrier for their support in establishing the website for the Data Contamination Database hosted on Hugging Face Spaces. We acknowledge the support of project Disargue (TED2021-130810B-C21, funded by MCIN/AEI /10.13039/501100011033 and by European Union NextGenerationEU/ PRTR) and the Basque Government (Research group funding IT-1805-22).

### v2

@@
-CONDA 2024 Shared Task on Evidence of Data Contamination: First Compilation Report
+Data Contamination Report from the 2024 CONDA Shared Task
 Abstract
@@
 Data Contamination Report from the 2024 CONDA Shared Task
-Oscar Sainz1 Iker García-Ferrero1 Alon Jacovi2 Jon Ander Campos3Yanai Elazar4,5Eneko Agirre1Yoav Goldberg2,4 Wei-Lin Chen6,7Jenny Chim8Leshem Choshen9,10Luca D’Amico-Wong11 Melissa Dell11Run-Ze Fan12Shahriar Golchin13Yucheng Li14Pengfei Liu12 Bhavish Pahwa15Ameya Prabhu17Suryansh Sharma18Emily Silcock11 Kateryna Solonko15David Stap19Mihai Surdeanu20Yu-Min Tseng21 Vishaal Udandarao16,17,22Zengzhi Wang12Ruijie Xu12Jinglin Yang11 1HiTZ Center - Ixa, University of the Basque Country UPV/EHU  2Bar Ilan University  3Cohere 4Allen Institute for Artificial Intelligence  5University of Washington  6National Taiwan University 7University of Virginia  8Queen Mary University of London  9MIT-IBM Watson AI Lab  10MIT 11Harvard University  12Shanghai Jiao Tong University  13University of Arizona  14University of Surrey 15Microsoft Research  16Tübingen AI Center  17University of Tübingen 18Indian Institute of Technology Kharagpur  19University of Amsterdam  20University of Arizona 21Microsoft  22University of Cambridge Contact: conda-workshop@googlegroups.com
+Oscar Sainz1 Iker García-Ferrero1 Alon Jacovi2 Jon Ander Campos3Yanai Elazar4,5Eneko Agirre1Yoav Goldberg2,4 Wei-Lin Chen6,7Jenny Chim8Leshem Choshen9,10Luca D’Amico-Wong11 Melissa Dell11Run-Ze Fan12Shahriar Golchin13Yucheng Li14Pengfei Liu12 Bhavish Pahwa15Ameya Prabhu16,17Suryansh Sharma18Emily Silcock11 Kateryna Solonko15David Stap19Mihai Surdeanu13Yu-Min Tseng6 Vishaal Udandarao17,21Zengzhi Wang12Ruijie Xu12Jinglin Yang11 1HiTZ Center - Ixa, University of the Basque Country UPV/EHU  2Bar Ilan University  3Cohere 4Allen Institute for Artificial Intelligence  5University of Washington  6National Taiwan University 7University of Virginia  8Queen Mary University of London  9MIT-IBM Watson AI Lab  10MIT 11Harvard University  12Shanghai Jiao Tong University  13University of Arizona  14University of Surrey 15Microsoft Research  16Tübingen AI Center  17University of Tübingen 18Indian Institute of Technology Kharagpur  19University of Amsterdam  20Microsoft  21University of Cambridge Contact: conda-workshop@googlegroups.com
 1 Introduction